Back to the NIKA run10 main page

Friday 7th November 2014

Barbara and Alessandro arrived at the telescope at 10h30. Alignment check. It was quite off (say 2-3cm on the input window). The spot isn't at the center of M5. I think that's OK. Also a few mm off (along y, a bit too high, in x it is OK) the center of M6, but I have adjusted M5 to shot on the center of the input window. Santiago played a bit with the compressed air controller but we decided it was good that way. We decided in any case to check again the alignement (but without touching if possible) on Tuesday before starting for real.

We have now a direct cable from the main NIKA console to the NIKA switch in the computers room. We spent some time playing for feeding the cable below the floor. Not so easy but the cable now is in place and it's comfortable.

Andrea in remote launched the MACs 35, 33 and 10. In local The acquisition seems OK. It's tuning. The configuration is that data are saved by archeops9 and MiniMac2 on Sami. Archeops9 saves data in local too.

Lost connection to the cryostat first, then to the electronics. In both cases the RJ45 connectors are really not very nice. Now it's OK but at any touch we risk the The strange 1-2-3 noise disappeared. I decide to regulate the temperature at 150mK as in run8.

The noise on the resonances seems quite OK at first view.

Barbara and Alessandro arrived at the telescope at 10h30. Alignment check. It was quite off (say 2-3cm on the input window). The spot isn't at the center of M5. I think that's OK. Also a few mm off (along y, a bit too high, in x it is OK) the ce

17h45. Raining. We wait.

We went very briefly to the sky for a couple of pointings. Just to see how off we are. Nicolas and Remi realised that there was a problem in the 1mm data in these scans. Despite horrible conditions the 2mm Uranus was OK and pretty close to the center of the array.

Saturday 8th November 2014

Nothing happened.

Sunday 9th November 2014

11h. We try to react to a GRB. Good excercise despite very low elevation (<20deg). Running to the cabin, turning the mirror and starting. We go to 3C279 fora focus liss. Found a FOCUS of -1.2mm while we where at -2mm. The quasar is easily seen on time traces at 2mm and 1mm.


Figure. Noise at 2mm taken during GRB observations (EL = 12 deg !!).

PUT LN2 in front of the window. Offsets are -250kHz and -450kHz. During the first slot (taumeter indicating roughly 0.15) we had up to -190KHz and -430KHz. So a bit better than LN2 but not that much. REMEMBER: zero has changed.

Monday 10th November 2014

9h. Sky is clear, wonderful day.

10-12h. We switch to NIKA to point GRB141109A. A few pointings and a focus on 1055+018 before. The pointing offsets are quite high this morning 18 and 3 arcsec. I set the pointing to that, focus to -1.6mm and I go to integrate on the GRB. A few glitches on some scans and only on a few pixels (3-4) of the 2mm. For the rest is good, observing more than one hour on source with tau=0.2 and at elevation of 30deg on average.

12h. Switch back to heterodyne.

A circular (GCN#17052, is issued by Alberto Castro-Tirado concerning GRB141109B that we observed last night during our technical time. Total telescope time on this GRB was last night roughly 1 hour.

With the window in front of the cryostat the resonances offsets wrt the param run10.txt are -380 and -850kHz respectively.

The technical time tonight is shifted to 19-23h.

During technical time OK again for pointing, focus etc. Small offsets of the order of 2-3 arcsec. Made a geometry and two skydips under poor sky conditions. Trying then to test the new pako scripts but it was difficult to manage. Then at 22h it started snowing so we stopped.

Tuesday 11th November 2014

10h. Snowing, wind, everything is stopped for today most probably.

Wednesday 12th November 2014

Observed in reasonable conditions between 20UT and 2UT (13/11). Projects Roberta Paladini (G114) and J-F Lestrade (debris disk)

Thursday 13th November 2014

During the night problems with tuning. In the morning too. Changed on mac10 the tuning_cmd.c (acqui) the parameter

#int sueil_angle_decalage 600

Was 400 mrad before.

A Lissajous scan on Uranus in a not very good weather:


Friday 14th November 2014

Alessandro went down to Granada, where a few tapas (no TAPAS!) will allow him to regain some energy. Martino took his place. Apparently the modification done by Alessandro worked. But we'll check more later: as of now everything is stopped (clouds, some rain, wind). Probably no observations until tomorrow, hopefully early in the morning.


Saturday 15th November 2014

In the morning, at around 8am, we get the hands back on the telescope: sky is clear! A bit of struggle to get things going (mirror was turned not in the good position, probably in the GISMO one... Plus, focus was terribly off due to deicing being stopped recently plus other important thermal deformations). When everything started to look ok, we reacted to a GRB that actually resulted to be just noise... Then, finally, 'science'!! A good share of SZ, then Roberta Paladini sources, now Jean Francois is taking control. A couple of hand made skydips as well. The first one in too good conditions so fit on 2mm pixels not very reliable..

Sunday 16th November 2014

Everything keeps going smoothly. Once the 'nikelB' box got blocked. But ok, recovered everything. Mac 35 keeps giving many errors, so sticking to mac9 as the reference one. Integrating on various project: although the tau is relatively poor (0.2/0.3), the atmosphere is very stable, so pretty good data arriving.

Monday 17th November 2014

Alexandre and Guilaine left, Fred arrived. Change also in the external observers (R Paladini and JF Lestrade out, M Albrecht and B Svoboda in). Again very sunny day (maybe, too sunny! -> too hot -> a bit to much humidity... but still less than when we were in the conditions shown a couple of days ago!). An error (this time of NickelA) was the good excuse to wake up a bit early. But ok, as we say in Italy, 'the morning has gold in its mouth'! Apart from that, again everything keeps going smoothly. A complete reboot of Mac 35 allowed it to start working properly again, so actually from the morning of 17/11 data in X35 dir should be ok. But since Mac 9 continues to do its job very well, let's stick with it! In the evening, moment of panic due to a test of some modification of the tuning process that resulted in a global chaos... Actually, we had kept a 'safe copy' of the soft but we had the bad chance of having at the same time hardware problems on one Nikel, so it was not straightforward to disentangle the problem. 3hrs of work needed to go back in operation mode.

Tuesday 18th November 2014

Yesterdays work did pay out in the end: now everything looks very stable, the tuning process is almost unbeatable (even going directly from 300K to the sky it goes smoothly!) and the skydip works perfectly! So it was worth the effort! Only strange thing: 2 pixels (not near in frequency, will check if they are near physically in the focal plane) show a bizarre noise jumping between 2 levels: see images below. This behaviour is irregular, and appears only from time to time. Then, it kinds of 'calm down'... Ideas? The involved pixel are, if I am correct, 9 and 20 (or, if I am wrong, 8 and 21..)


Wednesday 19th November 2014

Poor weather conditions in the morning getting slightly better (tau decreased from 0.5 to 0.25).

11:00 UTC : too much wind. Observations have been stopped.

c u later

16:00 UTC : still windy

hasta luego

[...But in the meantime, Alain took the chance to add a few extra improvements to 'acquisition'. Only on mac35, to keep the mac10 safe. Now, tuning with 35 and keeping an eye on it. The first impression is that everything is working as expected. But we will stay attentive for time being...]

18:47 UTC : back on the sky

Thursday 20th November 2014

7:30 UTC : too much wind. Observations have been stopped.

c u later

8:30 UTC : max wind speed = 30 m/s

Friday 21 November 2014

Still too much wind, up to 150 km/h during the night. Opacity up to 0.7. We will be leaving slighly earlier than expected. Let's hope the data we got will be enough.

DailyReports10 (last edited 2014-11-21 09:42:31 by NikaBolometer)